OpenStudy (anonymous):

1−secθ/tanθ+tanθ/1−secθ=−2cscθ Verify that each equation is an identity.

OpenStudy (anonymous):

mostly algebra i think, little if no trig

OpenStudy (anonymous):

if you put \(x=\cos(x)\) and \(b=\sin(x)\) then the left hand sides is \[\frac{1-\frac{1}{a}}{\frac{b}{a}}+\frac{\frac{b}{a}}{1-\frac{1}{a}}\]

OpenStudy (anonymous):

ok, I can see where you are going with this:)(:

OpenStudy (anonymous):

multiply the top and bottom of both parts sides by \(a\) and you get \[\frac{a-1}{b}+\frac{b}{a-1}\] then go ahead and add them

OpenStudy (anonymous):

ok! i understand! thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

like i said, almost no trig maybe on trig step at the end